• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 08:23:34 on June 3,2002

    Entry number 85141

    keyword=Shift Summary

    New Crew on Shift : Brian Milbrath,Pierre Moussiegt, Eric Voutier

    0:30 Call MCC for an update; we may be down for 2 more hours...

    2:15 Still no beam...

    2:30 iocha1 needs reboot, ask MCC.

    5:15 Beam back to the hall / move to the optics target for beam spot check
    Spot size was checked and found 105 x 74 on 1H03B

    5:40 Ask MCC to deliver high current. They get many many trips. The beam is

    .... Bad beamless shift !!